Jacke:

if cos 70° =0.342,give another angle whose cosine is 0.342

surjithayer:

cos x=cos (360-x) other angle =360-70=290

NotLeiland:

360-70=290***

NotLeiland:

@surjithayer wrote:
cos x=cos (360-x) other angle =360-70=290
How tf did I get 240

mikewwe13:

cos 70° = 0.342 To find another angle whose cosine is 0.342, we can use the inverse cosine function (cos⁻¹) on both sides of the equation: cos⁻¹(cos 70°) = cos⁻¹(0.342) 70° = 70.53° (rounded to two decimal places) Therefore, another angle whose cosine is 0.342 is approximately 70.53°.

Narad:

Another angle is \[-70º\]